How many sets of four numbers can be chosen from a collection of ten, without regard to order? For example, if I have a set of integers 1 through 10, a subset of four might be 1,2,3,and 4. If I count that as one possible subset, then I do not count 4,3,2,and 1, since I am not concerned with order. Thanks.
Answer by user_dude2008(1862)

10 C 4 = 10!/(4!(10-4)!) = 10!/(4!6!) = (10*9*8*7*6!)/(4!6!) = (10*9*8*7)/(4*3*2*1) = 210

Answer: 210 different possible combos